\begin{thebibliography}{0}

\bibitem{ChZ17}
S.~Chinaglia and S.~Zerbini, {A note on singular and non-singular black holes}.
\newblock \emph{Gen. Rel. Grav.} \textbf{49} (2017), 75.

\bibitem{Bar68}
J.~M. Bardeen, {Non-singular general-relativistic gravitational collapse}.
\newblock In \emph{{Proceedings of of International Conference GR5, Tbilisi,
  USSR}} (1968), page 174.

\bibitem{AyG98}
E.~Ayon-Beato and A.~Garcia, {Regular black hole in general relativity coupled
  to nonlinear electrodynamics}.
\newblock \emph{Phys. Rev. Lett.} \textbf{80} (1998), 5056--5059.

\bibitem{AyG99}
E.~Ayon-Beato and A.~Garcia, {New regular black hole solution from nonlinear
  electrodynamics}.
\newblock \emph{Phys. Lett.} \textbf{B464} (1999), 25.

\bibitem{AyG99b}
E.~Ayon-Beato and A.~Garcia, {Nonsingular charged black hole solution for
  nonlinear source}.
\newblock \emph{Gen. Rel. Grav.} \textbf{31} (1999), 629--633.

\bibitem{AyG00}
E.~Ayon-Beato and A.~Garcia, {The Bardeen model as a nonlinear magnetic
  monopole}.
\newblock \emph{Phys. Lett.} \textbf{B493} (2000), 149--152.

\bibitem{AyG05}
E.~Ayon-Beato and A.~Garcia, {Four parametric regular black hole solution}.
\newblock \emph{Gen. Rel. Grav.} \textbf{37} (2005), 635.

\bibitem{Ser16}
O.~Sert, {Regular black hole solutions of the non-minimally coupled $Y(R)F^2$
  gravity}.
\newblock \emph{J. Math. Phys.} \textbf{57} (2016) 032501.

\bibitem{Dym92}
I.~Dymnikova, {Vacuum nonsingular black hole}.
\newblock \emph{Gen.Rel.Grav.} \textbf{24} (1992), 235--242.

\bibitem{Dym02}
I.~Dymnikova, {Cosmological term as a source of mass}.
\newblock \emph{Class. Quant. Grav.} \textbf{19} (2002), 725--740.

\bibitem{Dym03}
I.~Dymnikova, {Spherically symmetric space-time with the regular de Sitter
  center}.
\newblock \emph{Int. J. Mod. Phys.} \textbf{D12} (2003), 1015--1034.

\bibitem{Hay06}
S.~A. Hayward, {Formation and evaporation of regular black holes}.
\newblock \emph{Phys. Rev. Lett.} \textbf{96} (2006), 031103.

\bibitem{BDD03}
K.~A. Bronnikov, A.~Dobosz and I.~G. Dymnikova, {Nonsingular vacuum cosmologies
  with a variable cosmological term}.
\newblock \emph{Class. Quant. Grav.} \textbf{20} (2003), 3797--3814.

\bibitem{BrF06}
K.~A. Bronnikov and J.~C. Fabris, {Regular phantom black holes}.
\newblock \emph{Phys. Rev. Lett.} \textbf{96} (2006), 251101.

\bibitem{BMD07}
K.~A. Bronnikov, V.~N. Melnikov and H.~Dehnen, {Regular black holes and black
  universes}.
\newblock \emph{Gen. Rel. Grav.} \textbf{39} (2007), 973--987.

\bibitem{BrD07}
K.~A. Bronnikov and I.~G. Dymnikova, {Regular homogeneous T-models with vacuum
  dark fluid}.
\newblock \emph{Class. Quant. Grav.} \textbf{24} (2007), 5803--5816.

\bibitem{Ans08}
S.~Ansoldi, {Spherical black holes with regular center: A Review of existing
  models including a recent realization with Gaussian sources}.
\newblock In \emph{{Conference on Black Holes and Naked Singularities Milan,
  Italy, May 10-12, 2007}} (2008).

\bibitem{AKM87}
A.~Aurilia, R.~S. Kissack, R.~B. Mann and E.~Spallucci, {Relativistic Bubble
  Dynamics: From Cosmic Inflation to Hadronic Bags}.
\newblock \emph{Phys. Rev.} \textbf{D35} (1987), 2961.

\bibitem{APS89}
A.~Aurilia, M.~Palmer and E.~Spallucci, {Evolution of Bubbles in a Vacuum}.
\newblock \emph{Phys. Rev.} \textbf{D40} (1989), 2511.

\bibitem{FMM89}
V.~P. Frolov, M.~A. Markov and V.~F. Mukhanov, {Through a black hole into a new Universe?}
\newblock \emph{Phys. Lett.} \textbf{B216} (1989), 272--276.
\newblock [,52(1990)].

\bibitem{FMM90}
V.~P. Frolov, M.~A. Markov and V.~F. Mukhanov, {Black Holes as Possible Sources
  of Closed and Semiclosed Worlds}.
\newblock \emph{Phys. Rev.} \textbf{D41} (1990), 383.

\bibitem{BaP90}
R.~Balbinot and E.~Poisson, {Stability of the Schwarzschild-de Sitter model}.
\newblock \emph{Phys. Rev.} \textbf{D41} (1990), 395--402.

\bibitem{NOS13}
P.~Nicolini, A.~Orlandi and E.~Spallucci, {The final stage of gravitationally
  collapsed thick matter layers}.
\newblock \emph{Adv. High Energy Phys.} \textbf{2013} (2013), 812084.

\bibitem{Nic05}
P.~Nicolini, {A model of radiating black hole in noncommutative geometry}.
\newblock \emph{J. Phys.} \textbf{A38} (2005), L631--L638.

\bibitem{NSS06a}
P.~Nicolini, A.~Smailagic and E.~Spallucci, {The fate of radiating black holes
  in noncommutative geometry}.
\newblock \emph{ESA Spec. Publ.} \textbf{637} (2006), 11.1.

\bibitem{NSS06b}
P.~Nicolini, A.~Smailagic and E.~Spallucci, {Noncommutative geometry inspired
  Schwarzschild black hole}.
\newblock \emph{Phys. Lett.} \textbf{B632} (2006), 547--551.

\bibitem{Riz06}
T.~G. Rizzo, {Noncommutative inspired black holes in extra dimensions}.
\newblock \emph{JHEP} \textbf{09} (2006), 021.

\bibitem{ANS07}
S.~Ansoldi, P.~Nicolini, A.~Smailagic and E.~Spallucci, {Noncommutative
  geometry inspired charged black holes}.
\newblock \emph{Phys. Lett.} \textbf{B645} (2007), 261--266.

\bibitem{SSN09}
E.~Spallucci, A.~Smailagic and P.~Nicolini, {Non-commutative geometry inspired
  higher-dimensional charged black holes}.
\newblock \emph{Phys. Lett.} \textbf{B670} (2009), 449--454.

\bibitem{Nic09}
P.~Nicolini, {Noncommutative black holes, the final appeal to quantum gravity:
  a review}.
\newblock \emph{Int. J. Mod. Phys.} \textbf{A24} (2009), 1229--1308.

\bibitem{NiS10}
P.~Nicolini and E.~Spallucci, {Noncommutative geometry inspired dirty black
  holes}.
\newblock \emph{Class. Quant. Grav.} \textbf{27} (2010), 015010.

\bibitem{SmS10}
A.~Smailagic and E.~Spallucci, {'Kerrr' black hole: the lord of the string}.
\newblock \emph{Phys. Lett.} \textbf{B688} (2010), 82--87.

\bibitem{MoN10b}
L.~Modesto and P.~Nicolini, {Charged rotating noncommutative black holes}.
\newblock \emph{Phys. Rev.} \textbf{D82} (2010), 104035.

\bibitem{Mod04}
L.~Modesto, {Disappearance of black hole singularity in quantum gravity}.
\newblock \emph{Phys. Rev.} \textbf{D70} (2004), 124009.

\bibitem{Mod06}
L.~Modesto, {Loop quantum black hole}.
\newblock \emph{Class. Quant. Grav.} \textbf{23} (2006), 5587--5602.

\bibitem{Mod08}
L.~Modesto, {Gravitational collapse in loop quantum gravity}.
\newblock \emph{Int. J. Theor. Phys.} \textbf{47} (2008), 357--373.

\bibitem{MMN11}
L.~Modesto, J.~W. Moffat and P.~Nicolini, {Black holes in an ultraviolet
  complete quantum gravity}.
\newblock \emph{Phys. Lett.} \textbf{B695} (2011), 397--400.

\bibitem{Nic12}
P.~Nicolini, {Nonlocal and generalized uncertainty principle black holes} {}
  arXiv:1202.2102 [hep-th].

\bibitem{BMM14}
C.~Bambi, D.~Malafarina and L.~Modesto, {Terminating black holes in
  asymptotically free quantum gravity}.
\newblock \emph{Eur. Phys. J.} \textbf{C74} (2014), 2767.

\bibitem{FZd15}
V.~P. Frolov, A.~Zelnikov and T.~de~Paula~Netto, {Spherical collapse of small
  masses in the ghost-free gravity}.
\newblock \emph{JHEP} \textbf{06} (2015), 107.

\bibitem{Fro15}
V.~P. Frolov, {Mass-gap for black hole formation in higher derivative and ghost
  free gravity}.
\newblock \emph{Phys. Rev. Lett.} \textbf{115} (2015), 051102.

\bibitem{FrZ16}
V.~P. Frolov and A.~Zelnikov, {Head-on collision of ultrarelativistic particles
  in ghost-free theories of gravity}.
\newblock \emph{Phys. Rev.} \textbf{D93} (2016), 064048.

\bibitem{FrZ17}
V.~P. Frolov and A.~Zelnikov, {Quantum radiation from an evaporating
  nonsingular black hole}.
\newblock \emph{Phys. Rev.} \textbf{D95} (2017) 124028.

\bibitem{flecture}
R.~P. Feynman, R.~B. Leighton and M.~Sands, \emph{The Feynman Lectures on
  Physics, Vol. II: The New Millennium Edition: Mainly Electromagnetism and
  Matter},  (Basic Books2011).

\bibitem{BaN93}
H.~Balasin and H.~Nachbagauer, {The energy-momentum tensor of a black hole, or
  what curves the Schwarzschild geometry? }.
\newblock \emph{Class. Quant. Grav.} \textbf{10} (1993), 2271.

\bibitem{BaN94}
H.~Balasin and H.~Nachbagauer, {Distributional energy momentum tensor of the
  Kerr-Newman space-time family}.
\newblock \emph{Class.Quant.Grav.} \textbf{11} (1994), 1453--1462.

\bibitem{BaN95}
H.~Balasin and H.~Nachbagauer, {The Ultrarelativistic Kerr geometry and its
  energy momentum tensor}.
\newblock \emph{Class. Quant. Grav.} \textbf{12} (1995), 707--714.

\bibitem{BaN97}
H.~Balasin, {Distributional energy momentum tensor of the extended Kerr
  geometry}.
\newblock \emph{Class. Quant. Grav.} \textbf{14} (1997), 3353--3362.

\end{thebibliography}

\documentclass{ws-ijgmmp}

\usepackage[utf8]{inputenc}

\begin{document}

\markboth{Authors' Names}
{Instructions for Typing Manuscripts (Paper's Title)}

%
\catchline{}{}{}{}{}
%

\title{REMARKS ON REGULAR BLACK HOLES}

\author{PIERO NICOLINI}

\address{Frankfurt Institute for Advanced Studies (FIAS), Ruth-Moufang-Str. 1\\Frankfurt am Main, D-60438, Germany\\
Institut f\"ur Theoretische Physik, Goethe-Universit\"at Frankfurt am Main, Max-von-Laue-Str. 1\\
Frankfurt am Main, D-60438, Germany\\
\email{nicolini@fias.uni-frankfurt.de} }

\author{ANAIS SMAILAGIC}

\address{Istituto Nazionale di Fisica Nucleare (INFN), Sezione di Trieste,
Strada Costiera 11\\
Trieste, I-34152, Italy\\
smailagic@ts.infn.it }

\author{EURO SPALLUCCI}

\address{Dipartimento di Fisica, Universit\`a degli Studi di Trieste \&
Istituto Nazionale di Fisica Nucleare (INFN), Sezione di Trieste,
Strada Costiera 11\\
Trieste, I-34152, Italy\\
spallucci@ts.infn.it}
\maketitle

\begin{history}
\received{(Day Month Year)}
\revised{(Day Month Year)}
\end{history}

\begin{abstract}
Recently it has been claimed by Chinaglia and Zerbini that the curvature singularity is present even in the so-called regular black hole solutions of the Einstein equations.  In this brief note we show that this criticism is devoid of any physical content.
\end{abstract}

\keywords{Exact solutions; Black holes; Space-time singularities, cosmic censorship, etc..}

\section{Introduction}
In a recent paper~\cite{ChZ17} Chinaglia and Zerbini (CZ) have claimed that a large number of non-singular black hole solutions of the Einstein equations still contain a curvature singularity,  \textit{e.g.} non-linear electrodynamics black holes~\cite{Bar68,AyG98,AyG99,AyG99b,AyG00,AyG05,Ser16}, vacuum filled spacetimes \cite{Dym92,Dym02,Dym03,Hay06}, phantom matter black holes~\cite{BDD03,BrF06,BMD07,BrD07}, regular gravitational shells~\cite{Ans08,AKM87,APS89,FMM89,FMM90,BaP90,NOS13}, noncommutative geometry inspired black holes~\cite{Nic05,NSS06a,NSS06b,Riz06,ANS07,SSN09,Nic09,NiS10,SmS10,MoN10b}, loop quantum black holes~\cite{Mod04,Mod06,Mod08} and
nonlocal gravity black holes~\cite{MMN11,Nic12,BMM14,FZd15,Fro15,FrZ16,FrZ17}.

The CZ reasoning goes as follows. The Einstein equations for a spherically symmetric, static, asymptotically
flat geometry reduce to the single ordinary differential equation: 
\begin{equation}
 r f^\prime + f = 1 -8\pi r^2 \rho \label{eff}
\end{equation}
where, $ g_{00}(r)=-f(r)=-1/g_{rr} $ and $\rho$ is the energy density. CZ claim that the general solution
of this equation is
\begin{equation}
\label{eq:zerbini}
f(r)=1-\frac{C}{r}-\frac{8\pi}{r}\int_0^r du \ u^2\rho(u)
\end{equation}
with $C$ an arbitrary constant. According to them, irrespective of the profile of $\rho$, the metric always exhibits a curvature singularity. This also holds for the  ``\emph{vacuum}'' case, \textit{i.e.},  $\rho\equiv 0$, whose solution would read
\begin{equation}
 f_0(r) = 1-\frac{C}{r}. \label{omo}
\end{equation}
%
Therefore, their claim is that all the papers 
dealing with regular black holes have implicitly assumed $C= 0$ without justification.

\section{Counter-arguments}
\subsection{``Vacuum solution'' and ``solution in vacuum'': massive rabbit out of the magician hat}
There is a very recurrent misunderstanding between ``vacuum solution'' and ``solution in vacuum''. To our knowledge,
 the clearest explanation is given by Richard Feynman in \cite{flecture} referring to  Poisson's equation in electrostatics:
 \begin{quotation}
 Suppose we want a solution of the equation in free space. The Laplacian is
zero because we are assuming that there are no charges anywhere. But the
spherically symmetric solution gives
\begin{equation}
\phi(r)= a + \frac{b}{r}
\end{equation}
in \emph{free} space. \emph{Something is evidently wrong}! The first constant term corresponds
to no charges (empty space) but we have a second term which says that there is a
point charge at the origin. So, although we are solving for a potential in free
space our solution gives the field for a point charge at the origin.
 \end{quotation}
The origin of the misunderstanding is due to the fact that solutions in vacuum are often obtained by solving the Laplace's equation. The mass or the charge (the rabbit) magically resurfaces by invoking the Gauss' theorem. This procedure is customarily used when the source of the Poisson's equation is no longer a function but a distribution.

The lack of distinction between ``vacuum'' and ``in vacuum'' has largely affected the Schwarzschild geometry that is often presented as a vacuum solution, namely
\begin{equation}
 R_{\mu\nu}=0\ ,\quad T_{\mu\nu}\equiv 0\ . \label{vuoto}
\end{equation}
Following the Feynman's reasoning, this conclusion is just wrong. In fact the Scharzschild metric is  a solution outside the source, \textit{i.e.,} in vacuum as 
shown in  \cite{BaN93,BaN94,BaN95,BaN97}.  In other words the Schwarzschild metric is obtained when
\begin{equation}
 \rho= \frac{M}{4\pi r^2}\delta(r) \label{singdens}
\end{equation}
describing a point-like mass $M$ particle at the origin. 
 Thus, following the Feynman's reasoning, the curvature  singularity in $r=0$ is nothing but a ``~warning~'' that something is wrong! The Schwarzschild solution is actually the sum of the ``~\emph{true}~'' 
homogeneous solution  $f=1$ and  the particular solution  $2M/r$ due to the presence of the Dirac-delta source.

The same considerations apply to the solution (\ref{omo}).  It is  \emph{not} a ``vacuum'' solution but a solution in vacuum, \textit{i.e.} it holds everywhere except in $r=0$. 

Apart from mathematical technicalities, the basic results of Einstein gravity is that curved geometries are induced by the presence of matter(energy). Claiming that a singular term is present even in the absence of matter is physically nonsensical.
The only  vacuum solution, compatible with asymptotic flatness, is Minkowski space-time.

\subsection{Simple counter-examples}
For extremely skeptical readers we give two physical counter-examples disproving the CZ line of thought.

\begin{enumerate}
 \item The charged ``shell''.  \\
We start with a general physics textbook example such as the charged empty sphere: 
\begin{equation}
  \nabla^2 \phi = 4\pi \rho,
 \end{equation}
where $\rho \propto \delta(r-R)$. According to CZ, the potential inside the shell would be given as 
\begin{equation}
 \phi= C_0 + \frac{C_1}{r}
\end{equation}
This is clearly nonsense  as there is no electrostatic field inside the shell as predicted by
the Gauss' law.

\item de Sitter metric.\\
Everybody knows that de Sitter space solves 
\begin{equation}
 R_{\mu\nu}=\Lambda g_{\mu\nu}
\end{equation}
On the contrary, following the CZ reasoning, one concludes that the de Sitter metric is
\begin{equation}
 f(r)=1 + \frac{C}{r} -\frac{\Lambda}{3} r^2
\end{equation}
and the curvature diverges in $r=0$. Conversely there is no need to impose $C=0$ since 
the singular term is just not present. 
\end{enumerate}

\section{Conclusions}
We conclude that the statements of the paper
 \cite{ChZ17} do not hold since they are conceptually wrong. CZ simply start from false premises and obtain incorrect conclusions. The false premise is that homogeneous solutions are solutions in vacuum, typically obtained in the case of the Schwarzschild geomety or  potentials generated by point like sources. The incorrect conclusion following such a false premise is that regular sources generate singular solutions. 
 
  As such the CZ work does not shed any light on the problem of regular black holes, nor disproves the existing solutions. 

\label{sec:concl}

\section*{Acknowledgments}

The work of P.N. has been supported by the project ``Evaporation of the microscopic black holes'' of the German Research 
Foundation (DFG) under the grant NI 1282/2-2.

\end{document}